------------------------------------------------------------ Date: Sat Feb 21 15:59:20 2004, Shift: 08:00-16:00 Supervisor: IGB, with: PJ Log book#15 pages 63 - 63 Shift Summary: ------------------------------------------------------------ 1. Good beam when we came. In next store, they had problem with Collimator. Result: Lowered lum. 2. Data: runs 10464 -> 10471 (still running), FS 4 deg B 230k MRS 40 deg B 320k 3. After a while (~9 am), Djamels problem repeated itself. After about an hour, we got the HV reset, as well as the anode ramp-up speed. The are as a default, by BRAHVO, set to 500. Flemming surgested only 100. Djamel probably didn't do this which might be why it tripped again. HV has been working fine since. Although D1 complained that it was too hot in the power supply. It was fixed though.What a shift! ------------------------------------------------------------ _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Sat Feb 21 15:59:35 2004
This archive was generated by hypermail 2.1.8 : Sat Feb 21 2004 - 15:59:57 EST